OpenStudy (ineedhelp10):

The diameter of a circle extends from (-15,2) to (3,2). If the area of this circle is the same as a square, what is a side length of that square?

OpenStudy (ineedhelp10):

@Directrix if you dont mind /.\

Directrix (directrix):

@ineedhelp10 Get the distance from (-15,2) to (3,2). The points lie on the same horizontal line so you can count units to get the length. Post what you get, okay?

OpenStudy (ineedhelp10):

I got 17

OpenStudy (ineedhelp10):

I mean 18

Directrix (directrix):

If 18 is the diameter of the circle, then the radius of the circle = 9.

Directrix (directrix):

Now, we have to get the area of the circle. A = pi * r^2 A = 3.14 * 9 * 9 A = ? @ineedhelp10

OpenStudy (ineedhelp10):

A=81pi @Directrix

Directrix (directrix):

Sorry. I don't think my notification system is working on OpenStudy. -------------------

Directrix (directrix):

Take this --> 81pi and approximate for pi. Multiply 81 times 3.14. Post what you get.

OpenStudy (ineedhelp10):

I got 254.46 is that what you get?

Directrix (directrix):

I had 254.34. I think that is close enough. Let me check.

Directrix (directrix):

Yours is correct.

OpenStudy (ineedhelp10):

How do you plug it in your calculator? I think i might be plugging it wrong then

Directrix (directrix):

area of this circle is the same as a square Then, 254.46 = x^2 where x is the side of the square. Recall that the area of a square is side times side. @ineedhelp10 254.46 = x^2 --> What the square root of 254.46? Post what you get.

OpenStudy (ineedhelp10):

15.95 @Directrix

Directrix (directrix):

Right again.

Directrix (directrix):

So, 15.95 is the side of the square. I don't know if you are supposed to round that. Check your directions and answer options.

OpenStudy (ineedhelp10):

it doesnt mention anything about rounding

Directrix (directrix):

I would go with the 15.95, then.

OpenStudy (ineedhelp10):

and thats the final answer then?

Directrix (directrix):

Well, here is the final question: what is a side length of that square? And, 15.95 is what we figured to be the side of the square.

OpenStudy (ineedhelp10):

it'll be 15.95 as well then

OpenStudy (ineedhelp10):

@Directrix

Directrix (directrix):

@ineedhelp10